please make sure you have the right answer. You are given two converging lenses that have focal lengths of 63 mm and 26 mm.

(a) Show how the lenses should be arranged to form a compound microscope that has a tube length of 157 mm. State which lens to use as the objective, which lens to use as the eyepiece.

Use the 26 mm focal length lens as the eyepiece and the 63 mm focal length lens as the objective.

Use the 63 mm focal length lens as the eyepiece and the 26 mm focal length lens as the objective.    


State how far apart to place the lenses.
........ cm

State what overall magnification you expect to get, assuming the user has a near point of 25 cm.
.........
Answer #1

we know that in compound microscope we have two lenses one is onjective lens , eye piece


objectve lens has less focal length , forms real and inveted image ,
eye piece has large focal length , works as magnifying glass.


form ula for magnification is
      


       M = (L/f0)(1+d/fe)


L is length of the tube , f0 objective lens focal length = 26 mm = 0.26 cm
           fe focal length of eye piece fe = 63 mm = 0.63 cm
           L = 157 mm = 15.7 cm

magnification M = (15.7/0.26)(1+25/0.63)= 2456.6